40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 266 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  3123   Sat Jun 26 05:02:04 2010 ranaHowToSVNSVN woes

Quote:

I am trying to get an actual complete install of the 40m svn on my machine. It keeps stopping at the same point:

 I have always seen this when checking out the 40m medm SVN on a non-Linux box. I don't know what it is, but Yoichi and I investigated it at some point and couldn't reproduce it on CentOS. I think its some weirdness in the permissions of tmp files. It can probably be fixed by doing some clever checkin from the control room.

Even worse is that it looks like the whole 'SVN' mantra has been violated in the medm directory by the 'newCDS' team. It could be that Joe has decided to make the 40m a part of the official CDS SVN, which is OK, but will take some retraining on our part.

  3122   Fri Jun 25 20:32:30 2010 kiwamuUpdateGreen Lockinggreen power on the PSL table

The power of the green beam generated on the PSL table should be about 650uW in terms of the shot noise.

       One of the important parameters we should know is the power of the green beam on the PSL table because it determines the SNR.

The green beam finally goes to a photo detector together with another green beam coming from the arm cavity, and they make a beat signal and also shot noise.

So in order to obtain a good SNR toward the shot noise at the photo detector, we have to optimize the powers.

If we assume the green power from the arm is about 650uW,  a reasonable SNR can be achieved when these powers are at the same level. 

To get such power on the PSL table, a 90% partial reflector is needed for picking it off from the PSL as we expected.

 


  power dependency of SNR

      Suppose two lasers are going to a photo detector while they are beating (interfering).

The beat signal is roughly expressed by

      [signal]  ~ E1EE1 E2*,

                     ~ 2 ( PP2)½ cos (phi), 

 where  E1 and Erepresent the complex fileds,  Pand Prepresent their powers and phi is a phase difference.

This equation tells us that the strength of the signal is proportional to  ( PP2)½  .

At the same time we will also have the shot noise whose noise level depends on the inverse square route of the total power;

          [noise] ~ ( PP2)½.

 According to the equations above, SNR is expressed by

        SNR = [signal] / [noise] ~ ( PP2)½  / ( PP2)½.

If we assume Pis fixed,  the maximum SNR can be achieved when  P2 goes to the infinity. But this is practically impossible.

Now let's see how the SNR grows up as the power P increases. There are two kinds of the growing phase.

    (1) When PP1 , SNR is efficiently improved with the speed of  P2½.

    (2) But  when P>   P1 , the speed of growing up becomes very slow. In this regime increasing of  P2 is highly inefficient for improvement of the SNR.

Thus practically PP is a good condition for the SNR.

At this point the SNR already reaches about 0.7 times of the maximum, it's reasonably good.

 


 power estimation

         According to the fact above, we just adjust the green powers to have the same power levels on the PSL table.

 The table below shows some parameters I assume when calculating the powers.

ITM transmissivity @ 532nm  Ti 1.5 %
ETM transmissivity @ 532nm Te 4.5 %
Transmissivity of the arm cavity @ 532nm T_cav 74.4 %
Transmissivity of the BS @ 532nm T_BS 97 %
Transmissivity of  PR1 and SR1 @ 532nm T_PR 90%
Transmissivity of the PMC @ 1064 nm T_pmc 65 %
The power of the green beam at the end station P_end 1 mW
The power of the PSL  P_psl 2 W
Conversion efficiency of the PPKTP eta 3 %/W

         Attached figure shows a simplified schematic of the optical layout with some numbers. 

By using those parameters we can find that the green beam from the arm cavity is reduced to 650uW when it reaches the PSL table.

To create the green beam with the same power level on the table, the power of 1064 nm going to the doubling crystal should be about 150mW.

This amount of the power will be provided by putting a 90% partial reflector after the PMC.

 

Attachment 1: optical_power.png
optical_power.png
  3121   Fri Jun 25 15:22:45 2010 JenneOmnistructureSAFETYNo entry to the 40m LVEA until further notice!

                                                                                                                                                                                                                                                                               

The 40m corner station crane is out of order, and it's stuck in a way that prohibits entry to the 40m LVEA / IFO room for safety.  The crane has been locked out / tagged out. 

Until further notice, absolutely no one may enter the 40m LVEA.  Work is permitted in the desk / control room areas.

Signs have been posted on all doors into the LVEA.  Please consider those doors locked out / tagged out.

                                                                                                                                                                                                                                                                               

  3120   Fri Jun 25 12:09:27 2010 kiwamuUpdateComputersGPIB controller of HP8591E

I've just stolen a GPIB controller, an yellow small box, from the spectrum analyzer HP8591E.

The controller is going to be used for driving the old spectrum analyzer HP3563A for a while.

Gopal and I will be developing and testing a GPIB program code for HP3563A via the controller.

Once after we get a new GPIB controller, it will be back to the original place, i.e. HP8591E.

 

--- GPIB controller ----

name: teofila

address: 131.215.113.106

  3119   Fri Jun 25 08:10:23 2010 josephbUpdateCDSDaily Downs Update

Yesterday afternoon I went to downs and acquired the following materials:

2 100 ft long blue fibers, for use with the timing system.  These need to be run from the timing switch in 1Y5/1Y6 area to the ends.

3 ADCs (PMC66-16AI6455A-64-50M) and 2 DACs (PMC66-16AO16-16-F0-OF), bringing our total of each to 8.

7 ADC adapter boards which go in the backs of the IO chassis, bringing our total for those (1 for each ADC) to 8.

There were no DAC adapter boards of the new style available.  Jay asked Todd to build those in the next day or two (this was on Thursday), so hopefully by Monday we will have those.

Jay pointed out there are different styles of the Blue and Gold adapter boxes (for ADCs to DB44/37) for example.  I'm re-examining the drawings of the system (although some drawings were never revised to the new system, so I'm trying to interpolate from the current system in some cases), to determine what adapter style and numbers we need.  In any case, those do not appear to have been finished yet (there basically stuffed boards in a bag in Jay's office which need to be put into the actual boxes with face plates).

When I asked Rolf if I could take my remaining IO chassis, there was some back and forth between him and Jay about numbers they have and need for their test stands, and having some more built.  He needs some, Jay needs some, and the 40m still needs 3.  Some more are being built.  Apparently when those are finished, I'll either get those, or the ones that were built for the 40m and are currently in test stands.

 

Edit:

Aparently Friday afternoon (when we were all at Journal Club), Todd dropped off the 7 DAC adapter boards, so we have a full set of those.

Things still needed:

1) 3 IO chassis (2 Dolphin style for the LSC and IO, and 1 more small style for the South end station (new X)).  We already have the East end station (new Y) and SUS chassis.

2) 2 50+ meter Ethernet cables and a router for the DAQ system.  The Ethernet cables are to go from the end stations to 1Y5-ish, where the DAQ router will be located.

3) I still need to finish understanding the old drawings drawings to figure out what blue and gold adapter boxes are needed.  At most 6 ADC, 3 DAC are necessary but it may be less, and the styles need to be determined.

4) 1 more computer for the South end station.  If we're using Megatron as the new IO chassis, then we're set on computers.  If we're not using Megatron in the new CDS system, then we'll need a IO computer as well.  The answer to this tends to depend on if you ask Jay or Rolf.

 

  3118   Fri Jun 25 01:28:33 2010 DmassHowToSVNSVN woes

I am trying to get an actual complete install of the 40m svn on my machine. It keeps stopping at the same point:

I do a

svn checkout --username svn40m https://nodus.ligo.caltech.edu:30889/svn /Users/dmass/svn

A blah blah blah many files

...

A    /Users/dmass/svn/trunk/medm/c1/lsc/C1LSC_ComMode.adl.28oct06
svn: In directory '/Users/dmass/svn/trunk/medm/c1/lsc'
svn: Can't copy '/Users/dmass/svn/trunk/medm/c1/lsc/.svn/tmp/text-base/C1LSC_MENU.adl.svn-base' to '/Users/dmass/svn/trunk/medm/c1/lsc/.svn/tmp/C1LSC_MENU.adl.tmp.tmp': No such file or directory

I believe I have always had this error come up when trying to do a full svn install. Any illumination is welcome.

 

 

  3117   Thu Jun 24 18:47:26 2010 FrankDAQIOOVME crate rebooted

we had to reboot the IOO VME crate right before lunch as the DAQ wasn't working correct meaning showing no real signals anymore, only strange noise. The framebuilder and everything else was working fine at that time.

  • The channel used for the phase noise measurement stopped showing any useful signal right after midnight, so all the other IOO-MC signals.
  • The data taken with those channels showed something like a 140 counts or so of steady offset with something which looked like the last bit fluctuating.
  • Whatever signal we connected to the input it didn't change at all, floating/shorted input, sine wave etc.
  • the other channels for the MC which we checked showed the same strange behaviour

As the other channels showed the same effect we decided to reboot the crate and everything was fine afterwards.

  3116   Thu Jun 24 16:59:24 2010 josephbUpdateVACFinished restoring the access connector and door

[Jenne,  Kiwamu, Steve, Sharmila, Katherine, Joe]

We finished bolting the door on the new ITMX (old ITMY) and putting the access connector section back into place.  We finished with torquing all the bolts to 40 foot-pounds.

  3115   Thu Jun 24 13:02:59 2010 JenneUpdateComputersSome lunchtime reboots

[Jenne, Megan, Frank]

We rebooted c1iovme, c1susvme1, and c1susvme2 during lunch.  Frank is going to write a thrilling elog about why c1iovme needed some attention.

C1susvme 1&2 have had their overflow numbers on the DAQ_RFMnetwork screen red at 16384 for the past few days.  While we were booting computers anyway, we booted the suses.  Unfortunately, they're still red.  I'm too hungry right now to deal with it....more to follow.

  3114   Thu Jun 24 11:16:32 2010 Sharmila, Rana and KiwamuHowToVACInspection of the BS (sorry, no sounds)
  3113   Thu Jun 24 06:49:29 2010 AidanUpdateGreen Lockinga channel for PPKTP temperature

Is this a setpoint temperature that we can change by writing to the channel or is it a readout of the actual temperature of the oven?

kiwamu:

This is a readout channel just to monitor the actual temperature.

Quote:

We added a channel on c1psl in order to monitor the temperature of the PPKTP sitting on the PSL table.

  3112   Thu Jun 24 01:02:34 2010 Sharmilla, Rana and KiwamuUpdateGreen Lockinga channel for PPKTP temperature

We added a channel on c1psl in order to monitor the temperature of the PPKTP sitting on the PSL table.

To take continuous data of the temperature we added the channel by editing the file: target/c1psl/c1psl.db

We named the channel "C1:PSL-PPKTP_TEMP".

To reflect this change we physically rebooted c1psl by keying the crate.

  3111   Wed Jun 23 23:55:03 2010 Katharine, Sharmila, Rana, Steve and KiwamuUpdateVACwiped the BS

Some unused optics in the BS chamber were removed. 

After that the beam splitter has been drag wiped. 

So now the BS chamber is waiting for the installation of  the other core optics i.e. PRM, SRM and Tip-Tilts. 

 

-- removing of unused optics

      There were some unused optics, mainly 1.5 inch optics which had been used for the oplevs in the chamber. 

Kathaine, Shamila (Team Magnet) and Kiwamu took those optics out from the chamber.

And then we carefully wrapped each of them by aluminum foils and put them in some clear boxes.

In fact, before wrapping them, we gently attached lens papers on their HR surfaces such that aluminum foils can not damage it.

Now there are only three 1.5 inch optics in the chamber, and they are supposed to be used for the oplevs.

We didn't remove any of the 2 inch optics and the PZT mirrors because they are still going to be used.

These are the pictures of the BS chamber after we cleaned up them. 

 

-- wiping of the BS

        Rana and Kiwamu drag wiped the HR surface of the BS by using lens paper with the solvents.

The below is the procedure we did. You can find some details about the wiping technique for suspended optics in this entry.

In this time we could wipe the beam splitter without removing the front earthquake stops because the beam splitter was brought close enough to us. 

 

(1). put some blocks attaching the edge of the bottom plate of the tower in order to record the original position.

(2).  locked the beam splitter to the frame by screwing the earthquake stops.

(3). made sure if it is really locked by seeing the output signal of the OSEMs in dataviewer. If it's locked successfully, the resonant frequency gets higher and the Q-value gets lower.

(3).  moved the BS tower close to the door in order to reach the beam splitter easily.

(4). inspected the surface by using a fiber light. There were about 10 bright spots on the HR surface.

(5). wiped the surface three times by using the lens paper with Aceton.

(6). wiped it several times with Isopropyl.

(7). inspected the surface again, found there were no big bright spots near the center. Thumbed up 

(8). put the tower back to the original place and released the beam splitter from the earthquake stops.

  3110   Wed Jun 23 23:08:30 2010 ranaConfigurationPSLFSS SLOWDC should be ~-4.0

 

  3109   Wed Jun 23 18:05:00 2010 KojiConfigurationPSLFSS SLOWDC should be ~-4.0

FSS SLOWDC slider is at around 0.

Please someone relock this at ~-4.0 to exploit some last juice of the fruit.

See this entry for the details of the operating point.

 

Attachment 1: C1PSL_FSS.png
C1PSL_FSS.png
  3108   Wed Jun 23 17:48:16 2010 steveUpdateMOPAlaser head temp

The laser chiller temp is fluctuating and the power output is decreasing. See 120 days plot.

Yesterday I removed ~300cc water from the overflowing chiller tank.

Attachment 1: htemp120d.jpg
htemp120d.jpg
  3107   Wed Jun 23 15:33:42 2010 josephbUpdateCDSDaily Downs Update

I visited downs and announced that I would be showing up again until all the 40m hardware is delivered. 

I brought over 4 ADC boards and 5 DAC boards which slot into the IO chassis.

The DACs are General Standards Corporation, PMC66-16AO16-16-F0-OF, PCIe4-PMC-0 adapters.

The ADCs are General Standards Corporation, PMC66-16AI6455A-64-50M, PCIe4-PMC-0 adapters.

These new ones have been placed with the blue and gold adapter boards, under the table behind the 1Y4-1Y5 racks.

With the 1 ADC and 1 DAC we already have, we now have enough to populated the two ends and the SUS IO chassis.  We have sufficient Binary Output boards for the entire 40m setup.  I'm going back with a full itemized list of our current equipment, and bring back the remainder of the ADC/DAC boards we're due.  Apparently the ones which were bought for us are currently sitting in a test stand, so the ones I took today were from a different project, but they'll move the test stand ones to that project eventually.

I'm attempting to push them to finish testing the IO chassis and the remainder of those delivered as well.

I'd like to try setting up the SUS IO chassis and the related computer this week since we now have sufficient parts for it.  I'd also like to move megatron to 1Y3, to free up space to place the correct computer and IO chassis where its currently residing.

  3106   Wed Jun 23 15:15:53 2010 josephbSummaryComputers40m computer security issue from last night and this morning

The following is not 100% accurate, but represents my understanding of the events currently.  I'm trying to get a full description from Christian and will hopefully be able to update this information later today.

 

Last night around 7:30 pm, Caltech detected evidence of computer virus located behind a linksys router with mac address matching our NAT router, and at the IP 131.215.114.177.  We did not initially recognize the mac address as the routers because the labeled mac address was off by a digit, so we were looking for another old router for awhile.  In addition, pings to 131.215.114.177 were not working from inside or outside of the martian network, but the router was clearly working.  

However, about 5 minutes after Christian and Mike left, I found I could ping the address.  When I placed the address into a web browser, the address brought us to the control interface for our NAT router (but only from the martian side, from the outside world it wasn't possible to reach it).

They turned logging on the router (which had been off by default) and started monitoring the traffic for a short time.  Some unusual IP addresses showed up, and Mike said something about someone trying to IP spoof warning coming up.  Something about a file sharing port showing up was briefly mentioned as well.

The outside IP address was changed to 131.215.115.189 and dhcp which apparently was on, was turned off.  The password was changed and is in the usual place we keep router passwords.

Update: Christian said Mike has written up a security report and that he'll talk to him tomorrow and forward the relevant information to me.  He notes there is possibly an infected laptop/workstation still at large.  This could also be a personal laptop that was accidently connected to the martian network.  Since it was found to be set to dhcp, its possible a laptop was connected to the wrong side and the user might not have realized this.

 

  3105   Wed Jun 23 12:52:35 2010 kiwamuUpdatePhotosBS chamber before cleaning up

  3104   Wed Jun 23 12:47:43 2010 JenneUpdateSUSPRM, SRM ready for vacuum

I fitzed with the PRM and SRM briefly, and I now believe that they're both ready to go into the chambers. 

For each optic, I used the microscope on a micrometer to check that the scribe lines on each side of the optic were at the same height.  Basic procedure was to center the microscope on one scribe line, move the microscope to the other side, to see how far the line was from center, and try to (very gently!!) rotate the optic in the wire about the z-axis about half the distance that the one scribe line needed to be.  Rinse and repeat several times until satisfied. 

I then checked that our HeNe oplev was still at 5.5" beam height, and that the beam traveled straight across the table.  I put the SRM in the oplev, unclamped the EQ stops, and waited for it to settle.  The HEPA filters were turned off, to minimize the breeze.  While the SRM settled, I worked on the height/rotation for the PRM on the other table. 

After checking the SRM balance, I clamped it and moved it, and checked the PRM balance, then turned off the HeNe and rewrapped everything in foil, and turned on the HEPAs.

Both the SRM and the PRM seem a little off in Pitch.  The beam returning to the QPD (placed just next to the laser) was always ~1cm above the center of the QPD.  The beam travel distance was ~3m (vaguely) from laser to optic to QPD.  This effect may be because the optics were originally balanced with OSEMs in place, and I didn't have any OSEMs today.  Koji and I found several months ago that the OSEMs have some DC affect on the optics.

Anyhow, since our optics are so small, I think the OSEMs and coils can handle this small DC offset in pitch, so I think we're ready to rock-n-roll with putting them in the chambers.

Still on the to-do list......Tip Tilts!

The photo shows the oplev beam position on (kind of) the QPD, for the SRM.  The PRM was basically the same.

Attachment 1: SRM_pitch_offset_small.jpg
SRM_pitch_offset_small.jpg
  3103   Wed Jun 23 12:31:36 2010 GopalUpdateGeneral6.16.10-6.23.10 Weekly Update

Summary of This Week's Activities:

6/16: LIGO Orientation; First Weekly Meeting; 40m tour with Jenne; Removed WFS Box Upper Panel, Inserted Cable, Reinstalled panel

6/17: Read Chapter 1 of Control Systems Book; LIGO Safety Meeting; Koji's Talk about PDH Techniques, Fabry-Perot Cavities, and Sensing/Control; Meeting w/ Nancy and Koji

6/18: LIGO Talk Part II; Glossed over "LASERS" book; Read Control Systems Book Chapter 2; Literary Discussion Circle

6/21: Modecleaner Matrix Discussion with Nancy; Suggested Strategy: construct row-by-row with perturbations to each d.f. --> Leads to some questions on how to experimentally do this.

6/22: Learned Simulink; Learned some Terminal from Joe and Jenne; LIGO Meeting; Rana's Talk; Christian's Talk; Simulink Intro Tutorial

6/23 (morning): Simulink Controls Tutorial; Successfully got a preliminary feedback loop working (hooray for small accomplishments!)

 

Outlook for the Upcoming Week:

Tutorials (in order of priority): Finish Simulink Tutorials, Work through COMSOL Tutorials

Reading (in order of priority): Jenne's SURF Paper, Controls Book, COMSOL documentation, Lasers by Siegman.

Work: Primarily COMSOL-related and pre-discussed with Rana

  3102   Wed Jun 23 12:28:34 2010 RazibSummaryPhase CameraWeeekly Summary

This past week I have completed the following tasks:

 

1. Built a trigger and power box for the camera GC 750M (06058) and took some test images to see whether the trigger box really works. Result: It is doing fine!

2. Went over the setup that is already sitting on the table. Ref: Aidan's elog entry

3. Attended seminars and talks given by Alan, Jahms, Koji and Rana.

4. Attended the mandatory laser safety training by Peter.

 

Expected task for this week (could be more):

1. Work out analytical expressions of the power of the carrier and sidebands going to the camera in the setup. (As suggested by Rana and Joe)

2. Work on producing beat signal to the camera using the He-Ne laser setup.

3. Move,if possible, to the Nd:YAG setup.

4. Go over the codes and paper by the past SURFers on the phase camera experiment.

 

trigger-box_circuit.png

 


 

Attachment 2: test1.png
test1.png
  3101   Wed Jun 23 11:31:12 2010 nancyUpdateWIKI-40M UpdateWeekly Update

This week I attended a whole lot of orientations, lectures, and meetings related to SURF. Done with general and laser safety training.

read Nergis' thesis for, and other material on WFS.

got confused with how the sidebands and shifted carrier frequencies are chosen for the Interferometer, read initial chapters of Regehr's thesis for teh same.

Made a plan for proceeding with the WFS work through discussions with Koji.

Understood the MC cavity and drew a diagram for it and the sensors.

Did Calculations for Electric field amplitudes inside and outside the MC cavity.

Saw the hardware of the WFS and QPD inside, and their routes to computers. Figured out which computer shows up the conditioned data from teh sensors.

Tried calculating the cavity axis for MC using geometry and ray tracing. Too complicated to be done manually.

Read some material (mainly Seigman) for physics of calculating the eigen-axis of the MC cavity with mirrors mis-aligned. Will calculate that using simulations, using the ABCD matrices approach.

Made a simple feedback simulink model yesterday to learn simulink. Made it run/compile. Saw the behaviour thru time signals at different points.

in the night, Made a simulink model of the sensor-mirror thing, with transfer functions for everything as dummy TFs. Compiles, shows signals in time. Remaining part is to put in real/near-real TFs in the model.

  3100   Wed Jun 23 11:25:14 2010 Katharine and SharmilaUpdateWIKI-40M UpdateMaglev

Weekly update


Lab work

We compared the magnetic field strength for 4 magnets in the original setup. The standard deviation was 3.15 G which corresponds to a variation of 2.4%. We had encountered difficulties with the stability of the Gaussmeter. The tip of the Gaussmeter was unsteady and wobbling which led to huge variations for a small change in distance. We stabilized the meter by taping it to a pencil and securing it with wire ties to an aluminum block. We then used translation stages to find the point of maximum field strength for each magnet, which allowed us much more stable readings.

Readings

We are reading and learning about feedback control systems. 

Modelling

Learning to model in Comsol. Our goals for the 1X1 model include incorporating the gravitational force in the measurements and find the distance for which attraction is the strongest, and experimenting with the mesh density and boundary conditions of the domain.

Meetings/seminars

Attended many meetings, including:
Laser safety training
SURF safety training
LIGO seminars
Journal club
LIGO experimental group meeting

  3099   Tue Jun 22 20:07:08 2010 JenneUpdate40m UpgradingFirst attempt at Tip Tilt hanging

[Jenne, Steve, Nancy, Gopal]

We made an attempt at hanging some of the Tip Tilt eddy current dampers today. 

Photo 1 shows the 2 ECDs suspended.

Procedure:

(1) Loosen the #4-40 screws on the side of the ECDs, so the wire can be threaded through the clamps.

(2) Place the ECDs in the locator jigs (not shown), and the locator jigs in the backplane (removed from main TT structure), all laying flat on the table.

(3) Get a length of Tungsten wire (0.007 inch OD = 180um OD), wipe it with acetone, and cut it into 4 ~8cm long segments (long enough to go from the top of the backplane to the bottom).

(4) Thread a length of wire through the clamps on the ECDs, one length going through both ECDs' clamps.

(5) One person hold the wire taught, and straight, and as horizontal as possible, the other person tightened the clamping screws on the ECDs.

(6) Again holding the wire in place, one person put the clamps onto the backplane (the horizontal 'sticks' with 3 screws in them).

(7) The end. In the future, we'll also clip off extra pieces of wire.

When we held up the backplane to check out our handy work, it was clear that the bottom ECD was a much softer pendulum than the top one, since the top one has the wire held above and below, while the bottom one only has the wire held on the top.  I assume we'll trim the wire so that the upper ECD is only held on the top as well?

Lessons learned:

* This may be a 3 person job, or a 2 people who are good at multitasking job.  The wire needs to be held, the ECDs need to be held in place so they don't move during the screwing/clamping process, and the screws need to be tightened.

* Make sure to actually hold the wire taught. This didn't end up happening successfully for the leftmost wire in the photo, and the wire is a bit loose between the 2 ECDs.  This will need to be redone.

* We aren't sure that we have the correct screws for the clamps holding the wire to the backplane.  We only have 3/16" screws, and we aren't getting very many threads into the aluminum of the backplane.  Rana is ordering some 316 Stainless Steel (low magnetism) 1/4" #4-40 screws.  We're going for Stainless because Brass (the screws in the photo), while they passed their RGA scan, aren't really good for the vacuum.  And titanium is very expensive.  

The 2nd photo is of the magnet sticking out of the optic holder.  The hole that the magnet is sitting in has an aluminum piece ~2/3 of the way through.  A steel disk has been placed on one side, and the magnet on the other.  By doing this, we don't need to do any press-fitting (which was a concern whether or not the magnets could withstand that procedure), and we don't need to do any epoxying.  We'll have to wait until the ECDs are hung, and the optic holder suspended, to see whether or not the magnet is sticking out far enough to get to the ECDs. 

Attachment 1: 2_ECDs_small.jpg
2_ECDs_small.jpg
Attachment 2: MagnetStickingOutFar_small.jpg
MagnetStickingOutFar_small.jpg
  3098   Tue Jun 22 18:56:32 2010 JenneUpdateEnvironmentBad placement of recycling bin

Someone has been moving the big blue recycling bin in front of the laser-chiller-chiller (the air conditioner in the control room).  This is unacceptable.  The chiller temp was up to 20.76C.  No good. 

You are free to move the recycling bin around so you can access drawers or the bike-exit-door in the control room, but make sure that it does not block air flow between the chiller-chiller and the chiller. 

The attached photo shows the BAD configuration.

Attachment 1: BlockingLaserChillerChiller_small.jpg
BlockingLaserChillerChiller_small.jpg
  3097   Tue Jun 22 13:38:13 2010 KojiUpdatePhase CameraCurrent phase camera setup. Seeing Acoustic beat

1. In terms of the AOM:

How much beam power is incident on the AOM? How much is the deflection efficiency?
i.e. How much is the power lost by the crystal, deflected in the 1st order, and remaining in the oth order?

I am curious because I assume the AOM (which vender?) is designed for 1064nm and the setup uses 632nm.

2. In terms of the EOM:

How much sidebands do you expect to have?

I assume the EOM is designed for 1064nm, the only difference is the coating at the end. Is this right? 

3. Beating

How much beating strength do you expect?

Is your beating level as expected?

How much is the contrast between the PM sideband and the frequency shifted carrier?
This must include the consideration on the presence of the carrier and the other sidebands.

Quote:

 We've set up a preliminary test bed for the phase camera. It simply uses a HeNe that is split into two beams. One is frequency shifted by an AOM by -40 MHz - df, where df is some acoustic frequency. The second beam is transmitted through a 40MHz EOM to get sidebands. The two beams are recombined and are, currently, incident on a photodetector, but this can be replaced by the phasecamera.

We turned everything on with df = 1kHz and confirmed that a 1kHz signal is visible on the output from the photodetector (PD). The signal looks to be about 1:300 of the DC level from the PD.

 

  3096   Tue Jun 22 09:45:21 2010 Aidan, Joe, RazibUpdatePhase CameraCurrent phase camera setup. Seeing Acoustic beat

 We've set up a preliminary test bed for the phase camera. It simply uses a HeNe that is split into two beams. One is frequency shifted by an AOM by -40 MHz - df, where df is some acoustic frequency. The second beam is transmitted through a 40MHz EOM to get sidebands. The two beams are recombined and are, currently, incident on a photodetector, but this can be replaced by the phasecamera.

We turned everything on with df = 1kHz and confirmed that a 1kHz signal is visible on the output from the photodetector (PD). The signal looks to be about 1:300 of the DC level from the PD.

Attachment 1: 2010-06-22_Phase_camera_layout_version_1.pdf
2010-06-22_Phase_camera_layout_version_1.pdf
  3095   Mon Jun 21 20:11:21 2010 KojiUpdatePhotosInspection of Magnets for the TTs

Were these magnets chipped before the Ni plating?

RA: Yes, it looks like this is the case. We also smashed some of the magnets against a metal surface and saw that a black grime was left. We should hold the magnets with a clean teflon clamp to measure the Gauss. Then we have to wipe the magnets before installing. I share Jenne's concern about the press-fit damaging the plating and so we need to consider using using glue or the ole magnetic attachment method. We should not rely on the structural integrity of the magnets at all.

  3094   Mon Jun 21 18:01:34 2010 JenneUpdate40m UpgradingSRM, PRM hung, magnets inspected

[Jenne, Kiwamu, Rana, Eric Gustafson]

The SRM and PRM have been re-hung, and are ready for installation into the chambers.  Once we put the OSEMs in, we may have to check the rotation about the Z-axis.  That was not confirmed today (which we could do with the microscope on micrometer, or by checking the centering of the magnets in the OSEMs).

Also, Eric and Rana inspected the Tip Tilt magnets, and took a few that they did their best to destroy, and they weren't able to chip the magnets.  There was concern that several of the magnets showed up with the coatings chipped all over the place.  However, since Rana and Eric did their worst, and didn't put any new chips in, we'll just use the ones that don't have chips in them.  Rana confiscated all the ones with obvious bad chips, so we'll check the strengths of the other magnets using a gaussmeter, and choose sets of 4 that are well matched. 

Eric, photographer extraordinaire, will send along the pictures he took, and we'll post them to Picasa.

  3093   Mon Jun 21 14:21:34 2010 Jenne, KiwamuUpdatePhotosInspection of Magnets for the TTs

Some pictures of "magnet inspection" from Picasa.

The coating of some magnets are chipped...

  3092   Sun Jun 20 18:28:25 2010 kiwamuUpdateGreen LockingRe: lock with PDH box

On the wiki I summarized about the modification of the PDH box which is currently running on the end PDH locking. 

http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/PDH_Universal_Box

The box was newly labeled "G1" standing for "Green locking #1".

Quote:

by using a modified PDH box the green laser on the X-end station is locked to the arm cavity.

  3091   Sun Jun 20 16:07:23 2010 KojiSummaryCOCCalibration of the metrology lab interferometer

Kiwamu and Koji

Summary

We have visited GariLynn's lab to make a calibration of the metrology interferometer. 

The newly calibrated value is

RoC(SRMU01) = 153.3+/- 1.6 [m]

This is to be compared with the specification of 142m +/- 5m

Although the calibration deviation from the previous value was found to be 1.3%, it is far from explaining the curvature difference between the spec (142m) and the measured value.


Motivation

The previous measurements of the SRM curvatures showed larger RoCs by ~10% compared with the spec.

It can be caused by the mis-calibration of the pixel size of the CCD in the metrology interferometer.
In order to confirm the calibration value, an object with known dimension should be measured by the instrument.

Method

We've got a flat blank optic from "Advanced Thin Film" together with a metalic ring.
The ring has been attached on the blank optic with 3 fragments of a double sided tape.
The RoC of SRMU1 was also measured in order to obtain "the radius of curvature of the day".

The calibration process is as follows:

  1. Measure the diameters of the ring by a caliper in advance to its attachment to the blank.
  2. Determine the inner and outer diameter of the ring in the obtained image.
    Note that the obtained image is pre-calibrated by the default value given by the measurement program
      (i.e. 0.270194mm/pixel for horizontal)
  3. Check the ratio of the diameters with the measured value by the caliper. Correct a systematic effect.
  4. Compare the image measurement and the caliper measurement.

Results

  1. The outer and inner diameters of 2.000" and 1.660" (measured by a caliper, error 0.005"). The ratio is 0.830+/-0.003.
  2. The center and radius for the inner circle were estimated to be (79.7924, 91.6372) and 21.4025 [mm].
    The center and radius for the outer circle were estimated to be (79.6532, 91.6816) and 25.6925 [mm].

    The error would be ~0.01mm considering they sweep 500 pixels by the circle and the pixel size is 0.27mm. i.e. 0.27/Sqrt(500) ~ 0.01mm
  3. Ratio of the inner and outer diameter is 0.8330 +/- 0.0005.
    The systematic error of x is given by solving (21.4025+x)/(25.6925-x)=0.83 ==> x = -0.042 +/- 0.043 [mm]. This is just a 0.2% correction.
    By correcting the above effect, we get (Rin, Rout) = (21.36 +/- 0.046, 25.74 +/- 0.047).
  4. By comparing the result with the caliper measurement, we get calibration factor of 1.013 +/- 0.005.
    This means we measured "1mm" as "1.013mm". The scale was too small.

    We have got the calibration of 0.2737+/-0.0014 [mm/pixel].

Discussion

Because of the calibration error, we measured too long RoC. The same day, we measured the curvature of SRMU01 as 155.26 m.
The newly calibrated value is

RoC(SRMU01) = 153.3+/- 1.6 [m]

This is the value to be compared with the specification of 142m +/- 5m

 

Attachment 1: ring1_inner_centering.pdf
ring1_inner_centering.pdf
Attachment 2: ring1_outer_centering.pdf
ring1_outer_centering.pdf
Attachment 3: SRMU01_pic.png
SRMU01_pic.png
  3090   Sat Jun 19 17:31:48 2010 ranaDAQCDSExcess Noise in C1:IOO-MC_DRUM1 fixed by reboot

I was getting an excess noise in the C1:IOO-MC_DRUM1 channel - it was a flat spectrum of 10 cts/rHz (corresponding to 600 uV/rHz).

I tried a few things, but eventually had to power cycle the crate with c1iovme in order to recover the standard ADC noise level of 3x10^-3 cts/rHz with a 1/sqrt(f) knee at 10 Hz.

I checked the gain of the channel by injecting a 2 Vpp sine wave at 137.035 Hz. 2Vpp as measured on a scope gives 31919 cts instead of the expected 32768, giving a 2.5% error from what we would have naively calculated.

Even so, the noise in this channel is very surprisingly good: 0.003 / (31919 / 2) = 187 nV /rHz.  The best noise I have previously seen from an ICS-110B channel is 800 nV/rHz. What's going on here?

  3089   Fri Jun 18 22:12:29 2010 AlbertoUpdateIOOMode Profile after Mode Matching Telescope (Round 2)

GJ

  3088   Fri Jun 18 21:45:39 2010 kiwamuUpdateIOOMode Profile after Mode Matching Telescope (Round 2)

           [Joe, Kiwamu]

The better mode overlap of 99.3% was achieved by moving MMT2 by ~5 cm 

In the past measurement (elog entry #3077) we already succeeded in getting 99.0% mode overlap.

But according to the calculation there still was a room to improve it by moving MMT2 by 10 cm.

Today we moved MMT2 by ~5 cm which is a reasonable amount we could move because of the narrow space in the chamber.

Eventually it successfully got the better mode overlap.

So we eventually finished mode matching of the new IOOs 


(details)

     Actually moving of MMT2 was done by flipping the mount without moving the pedestal post as Koji suggested. 

At the same time we also flipped the mirror itself (MMT2) so that the curved surface is correctly facing toward the incident beam.

By this trick, we could move the position of MMT2 without losing precious available space for the other optics in the OMC chamber.

     The attached plot shows the result of the mode measurement after the MMT.

During the fitting I neglected the data at x=27 m and 37 m because the beam at those points were almost clipped by the aperture of the beam scan.

- - Here are the fitting results

w0_v             = 2.81183       +/- 7.793e-03  mm  (0.2772%)

w0_h             = 2.9089        +/- 1.998e-02  mm  (0.687%)

z_v             = 5.35487        +/- 0.2244     m   (4.19%)

z_h             = 1.95931        +/- 0.4151     m   (21.18%)

All the distances are calibrated from the position of MMT2 i.e. the position of MMT2 is set to be zero.

        In order to confirm our results, by using the parameters listed above I performed the same calculation of mode overlaps as that posted on the last entry (see here)

The result is shown in Attachement 2. There is an optimum point at x=62mm.

This value is consistent with what we did because we moved MMT2 by ~5 cm instead of 10 cm. 

 

Attachment 1: MMT20100618_edit.png
MMT20100618_edit.png
Attachment 2: newIOO_overlap_edit.png
newIOO_overlap_edit.png
  3087   Fri Jun 18 15:07:26 2010 AlbertoUpdateLSCShort Cavity Length Adjustments

Quote:

You should have been in my lecture yesterday!
Power in the cavity is not a good index (=error signal) to judge the optimal length.
You should look at the phases of the length signals. (i.e. demodulation phase which gives you the maximum amplitude for CARM, PRC, SRC, etc)

You must move the SRC and PRC lengths at the same time.
The resonance of f1 (mostly) depends on the PRC length, but that of f2 depends on both the PRC and SRC lengths. 

Right. Ultimately the phase gain inside the cavity is what we look at. Calculating that for the SBs inside PRC and SRC is actually the first thing I did.

But I kept getting very small angles. Too small, I thought. Maybe there was some problem in the way I calculated it.

Then I made a power analysis to check if the SBs were getting affected at all by that 0.7degree phase shift they're picking up in the arms.

I wanted to show the point where I am, before leaving. But, I keep working on it.

  3086   Fri Jun 18 13:47:20 2010 KojiUpdateLSCShort Cavity Length Adjustments

You should have been in my lecture yesterday!
Power in the cavity is not a good index (=error signal) to judge the optimal length.
You should look at the phases of the length signals. (i.e. demodulation phase which gives you the maximum amplitude for CARM, PRC, SRC, etc)

You must move the SRC and PRC lengths at the same time.
The resonance of f1 (mostly) depends on the PRC length, but that of f2 depends on both the PRC and SRC lengths.

  3085   Fri Jun 18 13:42:52 2010 KojiHowToGeneralUpdate your work

All SURFs (and all others as always) are supposed to post the update of your status on the elog.

In fact, I already heard that Sharmila had been working on the serial connection to TC-200 and made some results. All of us like to hear the story.

  3084   Thu Jun 17 17:09:44 2010 AlbertoUpdateLSCShort Cavity Length Adjustments

I calculated the phase shifts that the sidebands would pick up in the arms in the case we changed the arm length to 38.4m as proposed. I obtained the following values (in degrees):

phi(-f2) = 0.66; phi(-f1) = -0.71; phi(f1) = 0.71; phi(+f2) = -0.66

These are the plots with the results as I obtained from an Optickle simulation (the second zooms in around 38.4m).

sidebandPhaseRotation_73430639654.png sidebandPhaseRotation_73430656054.png

These values agree with what Koji had already estimated (see elog entry 3023).

Since we can't make the arm longer than that, to increase the distance from the resonance, we would like to adjust the length of the short cavities to compensate for that.  For f2 (=55MHz), 0.7 degrees correspond to about 5cm. That is about the length change that we expect to make to the design.

I simulated with Optickle the effect of changing the length of either the SRC or the PRC. The best way I found to do that, was to measure the cavity circulating power when the macroscopic lengths change.

The following plots show the effect of changing either the PRC or SRC length (left or right figure), on the circulating power of both cavities at the same time (top and bottom plots).

shortCavityCirculatingPower_73430666992.png prcCirculatingPower_73430665955.png

 You can compare these with the case of perfect antiresonance as in the following plots:

shortCavityCirculatingPower_73430668892.png shortCavityCirculatingPower_73430669604.png

It seems that the design length for the short cavities are not too bad. f1 is not optimized in the PRC, but changing the length of the cavity wold just make f2 worse in SRC.

These simulations seem to support the choice of not changing the design cavity lengths for PRC and SRC.

Of course these are only an "open loop" simulations. At the moment we don't know what would be the effect of closing the control loops. That is something I'm going to do later. It'll be part of my studies on the effects of cavity absolute length on the whole IFO.

  3083   Wed Jun 16 18:44:07 2010 AlbertoConfigurationComputers40MARS

Quote:

i added my laptop's mac address to teh martian at port 13 today.

 

No personal laptop is allowed to the martian network. Only access to the General Computing Side is permitted.

Please disconnect it.

  3082   Wed Jun 16 18:14:13 2010 AidanUpdateGeneralGlass cover from overhead light smashed on PSL table

I was giving a tour of the 40m yesterday. We were looking at the PSL table. About 30 seconds after I turned the lights on a glass cover from one of the lights (NW corner) popped out of its holder and smashed on the table.

I've cleaned up all the broken glass I could see but there may be some small shards there. Please use caution in that area.

Attachment 1: DSC_1769.JPG
DSC_1769.JPG
Attachment 2: DSC_1768.JPG
DSC_1768.JPG
  3081   Wed Jun 16 18:12:16 2010 nancyConfigurationComputers40MARS

i added my laptop's mac address to teh martian at port 13 today.

 

  3080   Wed Jun 16 11:31:19 2010 josephbSummaryComputersRemoved scaling fonts from medm on Allegra

Because it was driving me crazy while working on the new medm screens for the simulated plant, I went and removed the aliased font entries in /usr/share/X11/fonts/misc/fonts.alias that are associated with medm.  Specifically I removed the lines  starting with widgetDM_.  I made a backup in the same directory called fonts.alias.bak with the old lines.

Medm now behaves the same on op440m, rosalba, and allegra - i.e. it can't find the widgetDM_ scalable fonts and defaults to a legible fixed font.

  3079   Tue Jun 15 21:28:44 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

I calculated the frequency of the double cavity pole for the 40m SRC-arm coupled cavity.

w_cc = (1 + r_srm)/(1- r_srm) * w_c

where w_c is the arm cavity pole angular frequency [w_c = w_fsr * (1-r_itm * r_etm)/sqrt(r_itm*r_etm) ]

I found the pole at about 160KHz. This number coincides with what I got earlier with my optickle model configured and tuned as I said in my previous entry. See attachments for plots of transfer functions with 0 and 10pm DARM offsets, respectively.

I think  the resonance at about 20 Hz that you can see in the case with non-zero DARM offset, is due to radiation pressure. Koji suggested that I could check the hypothesis by changing either the mirrors' masses or the input power to the interferometer. When I did it frequency and qualty factor of the resonance changed, as you would expect for a radiation pressure effect.

Independently, Jan also calculated the pole frequency of the transfer function DARM / ASQ2 as we would expect it for the SRC-coupled cavity. He also found the pole at about 160KHz. I'm attaching the plot with the transfer function he calculated.
He also said that the little bump at the pole frequency is OK considering that our signal recycling cavity is not much shorter than the arms.

This gave me more confidence about my optickle model of the 40m. This is quite comforting since I used that model other times in the past to calculate several things (i.e. effects of higher unwanted harmonics from the oscillator, or, recently, the power at the ports due to the SB resonating in the arms).

I don't know anymore what Valera said that wasn't right.
Also, as he said, he set it for the carrier to be resonant in the SRC-arms couple cavity. But that is not our case.
Attachment 1: allTransferFunctions_DARMoff_0.pdf
allTransferFunctions_DARMoff_0.pdf allTransferFunctions_DARMoff_0.pdf allTransferFunctions_DARMoff_0.pdf
Attachment 2: allTransferFunctions_DARM2AS_10pmDARMoffset.pdf
allTransferFunctions_DARM2AS_10pmDARMoffset.pdf allTransferFunctions_DARM2AS_10pmDARMoffset.pdf allTransferFunctions_DARM2AS_10pmDARMoffset.pdf
Attachment 3: Jan_DARM2AS.pdf
Jan_DARM2AS.pdf
  3078   Tue Jun 15 19:24:25 2010 ranaUpdateIOOMode Profile after Mode Matching Telescope

Quote:

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap.

 That's hot stuff.

  3077   Tue Jun 15 16:28:32 2010 kiwamuUpdateIOOMode Profile after Mode Matching Telescope

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap. 

Yesterday Jenne and I put MMT2 on the OMC table. MMT2 was carefully put by measuring the distance between MMT1 and MMT2.

The position looked almost the same as that drawn on the CAD design.

After putting it we measured the profile after the MMT.

 

The attached figure shows the computed mode overlap according to the fitting result while changing the position of MMT2 in a program code.

The x-axis is the position of MMT2, the current position is set to be zero. The y-axis is the mode match overlap.

Right now the overlap is 99.0% successfully, but this is not an optimum point because the maximum overlap can be achieved at x=100 mm in the plot.

It means we can have 99.6% by moving the position of MMT2 by another 10 cm. This corresponds to an expansion of the MMT length.

If this expansion is difficult due to the narrow available space in the chamber, maybe staying of MMT2 at the current position is fine.

Attachment 1: newIOO_overlap_edit.png
newIOO_overlap_edit.png
  3076   Mon Jun 14 22:16:08 2010 JenneUpdateIOOMode scan after Mode Matching Telescope

[Jenne, Kiwamu]

We measured the mode after the Mode Matching Telescope. 

 

---- fitted parameters ----

       w0_h =  2.85 +/- 0.0115 mm

       w0_v =  2.835 +/- 0.00600 mm

       z0_h = 5.4 +/- 0.447  m

       z0_v  = 6.9 +/- 0.305  m

Attachment 1: after_MMT.png
after_MMT.png
  3075   Mon Jun 14 07:57:07 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

  3074   Sun Jun 13 08:28:44 2010 valeraUpdateLocking40m Upgrade Optickle Model

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

ELOG V3.1.3-